# Incommensurate Twisted Bilayer Graphene: emerging quasi-periodicity and stability

### v1

@@ B.1 Diophantine condition from $\mathbb{R}$ to $\mathbb{R}^{2}$
and so, for $k\in\zeta$, because the maximal angle between $k$ and $\frac{\partial}{\partial\theta}(\frac{f^{\prime}}{|f^{\prime}|})$} is $\alpha+\frac{\pi}{4}$,
$|k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}}).$ (120)
@@
$|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}})-|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|.$ (121)
@@
$|k|\geq\eta+\frac{|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|}{|{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}})}\equiv R_{2}$ (122)
@@
$|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>\eta|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{4}}).$ (123)
@@
$|\Omega^{c}(x,f)|\leq\frac{2C_{3}}{\eta{\displaystyle\min_{\theta\in[\theta_{0},\theta_{1}]}}|\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})|\cos(\alpha+\frac{\pi}{4})}\sum_{k\in\mathbb{Z}^{2}\setminus\{0\}}|k|^{-1-\epsilon}.$ (124)
@@
B.2 Applying the Diophantine condition to $|p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}+lb+mb^{\prime}|$
We now apply lemma B.1 repeatedly to prove lemma IV.1.
@@
$|p_{F,i}^{\omega}-p_{F,j}^{\omega^{\prime}}+lb+mb^{\prime}|\equiv|lb+mb^{\prime}|\equiv|M|\geq\frac{C_{0}}{|m|^{\tau}}$ (125)
@@
Now, recalling the definition (99), we have that if $\theta\in\mathcal{D}(0,f_{\omega})$, then the inequality (28) with $i=i^{\prime}$, $\omega=\omega^{\prime}$, and $y=m$ holds with $C_{0}:=\frac{2\pi}{\sqrt{3}}C_{1}$. We therefore just need to use Lemma B.1 to ensure that the measure of this set is large. Taking $\theta_{0},\theta_{1}$ sufficiently small, it suffices to verify the conditions at $\theta=0$, and conclude by continuity. In particular, when $\theta_{0},\theta_{1}$ are small, $f^{\prime}(\theta)$ will take values in a small cone $\mathcal{C}_{f^{\prime}(0)}(\alpha)$ with $\alpha=O(\theta_{1})$. Next, by a straightforward computation, we find
$|f^{\prime}_{\omega}(0)|=\sqrt{\frac{5}{3}},\quad\left|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}(0)}{|f^{\prime}_{\omega}(0)|}\right|=\frac{2\sqrt{3}}{5}$ (128)
Which are both non-zero, so (103) is satisfied for small enough $\theta$. Since $x=0$, the other assumptions trivially hold: we choose $C_{3}<1/\sqrt{2}$ and $\eta<1$ such that $R_{1},R_{2}<1$, in which case the minima in (104) and (106) are taken over empty sets, so (104) and (106) hold trivially. Thus, by Lemma B.1, choosing $C_{1}=O(\theta_{1}^{2})$, the set $\mathcal{D}(0,f_{\omega})$ has a large measure.
We now repeat the argument for the other values of $i,j$, $y$, and $\omega,\omega^{\prime}$. First, note that $p_{F}^{+}-p_{F}^{-}=\frac{1}{3}(b_{1}-b_{2})$ so the condition (28) holds for $\omega\neq\omega^{\prime}$ whenever it holds for $\omega=\omega^{\prime}$. Next, note that
$|p_{F,i}^{\omega}-p_{F,j}^{\omega}+lb+mb^{\prime}|=|R^{T}(p_{F,i}^{\omega}-p_{F,j}^{\omega}+lb+mb^{\prime})|$ (129)
@@
$|p_{F,1}^{\omega}-p_{F,2}^{\omega}+lb+mb^{\prime}|\geq\frac{C_{0}}{|m|^{\tau}}$ (130)
@@
$|p_{F,1}^{\omega}-p_{F,2}^{\omega}+lb+mb^{\prime}|\geq\frac{2\pi}{\sqrt{3}}\left(x_{\omega}(\theta)+l_{1}+m\cdot f_{\omega}(\theta)\right)$ (131)
@@
Therefore, if $\theta\in\mathcal{D}(x_{\omega},f_{\omega})$, then (130) holds with $C_{0}=\frac{2\pi}{\sqrt{3}}C_{1}$. To show that this set has a large measure, we check the assumptions of Lemma B.1, as we did above. Again, we check the assumptions at $\theta=0$, and argue by continuity. We compute
$|x^{\prime}_{\omega}(0)|=\frac{1}{\sqrt{3}},\quad\left|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}(0)}{|f^{\prime}_{\omega}(0)|}\right|=\frac{8}{5\sqrt{15}}$ (133)
We thus find that if $C_{3}<1/\sqrt{2}-1/\sqrt{5}$ and $\eta<1-4\sqrt{2}/\sqrt{45}$, then $R_{1},R_{2}<1$, so the minima in (104) and (106) are taken over empty sets, so (104) and (106) hold trivially.
@@
All in all, we have found that if we restrict the values of $\theta$ to an intersection of Diophantine sets:
$\theta\in\bigcap_{\omega=\pm}\bigcap_{\sigma=\pm}\mathcal{D}(0,f_{\omega}(\sigma\theta))\cap\bigcap_{\omega=\pm}\bigcap_{\sigma=\pm}\mathcal{D}(x_{\omega}(\sigma\theta),f_{\omega}(\sigma\theta))$ (134)
@@
then (28) is satisfied for any value of $i,i^{\prime}$, $\omega,\omega^{\prime}$, and $y$ with a constant $C_{0}=O(\theta_{1}^{2})$. Because each set has an arbitrarily large measure (relative to $\theta_{0},\theta_{1}]$), their intersection also does.
@@ Appendix C Naive perturbation theory
The Schwinger function is computed using perturbation theory: formally,
@@ D.1 Symmetries of the system
$H_{1}$ and $H_{2}$ are invariant under (141) for the same reason as the conjugation and particle-hole symmetries. For the interlayer hopping, it is easiest to use the expression of $V$ in (II), which involves two integrals: one over $\widehat{\mathcal{L}}_{1}$ and one over $\widehat{\mathcal{L}}_{2}$. Let us discuss the invariance of the integral over $\widehat{\mathcal{L}}_{1}$, as the invariance of the other follows from a similar argument. We change variables in the integral: $k\mapsto-k+b_{1}+b_{2}$, as well as in the sum over $l$: $(l_{1},l_{2})\mapsto-(l_{1}+1,l_{2}+1)$, and in the sum over $\alpha$: $\alpha\mapsto\bar{\alpha}$. This changes $\tau_{l,\alpha}^{(1)}(k+lb)$ to $\tau_{-l-(1,1),\bar{\alpha}}^{(1)}(-k-lb)$. Now, by (13),

### v2

@@ B.1 Diophantine condition from $\mathbb{R}$ to $\mathbb{R}^{2}$
and so, for $k\in\zeta$, because the maximal angle between $k$ and $\frac{\partial}{\partial\theta}(\frac{f^{\prime}}{|f^{\prime}|})$} is $\alpha+\frac{\pi}{2}-\beta$,
$|k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{2}}-\beta).$ (124)
@@
$|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{2}}-\beta)-|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|.$ (125)
@@
$|k|\geq\eta+\frac{|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|}{|{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{2}}-\beta)}\equiv R_{2}$ (126)
@@
$|{\textstyle\frac{\partial}{\partial\theta}(\frac{x^{\prime}(\theta)}{|f^{\prime}(\theta)|})}+k\cdot{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|>\eta|k||{\textstyle\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})}|\cos(\alpha+{\textstyle\frac{\pi}{2}}-\beta).$ (127)
@@
$|\Omega^{c}(x,f)|\leq\frac{2C_{3}}{\eta{\displaystyle\min_{\theta\in[\theta_{0},\theta_{1}]}}|\frac{\partial}{\partial\theta}(\frac{f^{\prime}(\theta)}{|f^{\prime}(\theta)|})|\cos(\alpha+\frac{\pi}{2}-\beta)}\sum_{k\in\mathbb{Z}^{2}\setminus\{0\}}|k|^{-1-\epsilon}.$ (128)
@@
B.2 Applying the Diophantine condition to $|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|$
We now apply lemma B.1 repeatedly to prove lemma V.1. Throughout this section, we set $\beta=\frac{\pi}{4}$ and take $\alpha$ very small. Our proof is computer assisted.
@@
$|K^{i}_{\omega}-K^{j}_{\omega^{\prime}}+lb+mb^{\prime}|\equiv|lb+mb^{\prime}|\equiv|M|\geq\frac{C_{0}}{|m|^{\tau}}$ (129)
@@
Now, recalling the definition (103), we have that if $\theta\in\mathcal{D}(0,f_{\omega})$, then the inequality (4) with $i=i^{\prime}$, $\omega=\omega^{\prime}$, and $y=m$ holds with $C_{0}:=\frac{2\pi}{\sqrt{3}}C_{1}$. We therefore just need to use Lemma B.1 to ensure that the measure of this set is large. Taking $\theta_{0}$ and $\theta_{1}$ to be sufficiently close to each other (so that the value of $\alpha$ is arbitrarily small and the cone $\mathcal{C}_{\xi}(\alpha)$ is arbitrarily narrow), it suffices to verify the conditions at $\alpha=0$, and conclude by continuity.
We first verify (107). In Figure 4, we see that neither $f^{\prime}$ nor $\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}$ vanish for any value of $\theta\in[0,2\pi)$, so (107) holds.
Figure 4: $|f^{\prime}_{\omega}(\theta)|$ and $|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}|$ (blue: $\omega=+$, red: $\omega=$-, color online). These show (107) holds: neither of these quantities reaches 0.
Since $x=0$, the other assumptions trivially hold: we choose $C_{3}<1/\sqrt{2}$ and $\eta<1$ such that $R_{1},R_{2}<1$, in which case the minima in (108) and (110) are taken over empty sets, so (108) and (110) hold trivially. Thus, by Lemma B.1, choosing $C_{1}\ll(\theta_{1}-\theta_{0})^{2}$, the set $\mathcal{D}(0,f_{\omega})$ has a large measure.
We now repeat the argument for the other values of $i,j$, $y$, and $\omega,\omega^{\prime}$. First, note that $K_{+}-K_{-}=\frac{1}{3}(b_{1}-b_{2})$ so the condition (4) holds for $\omega\neq\omega^{\prime}$ whenever it holds for $\omega=\omega^{\prime}$. Next, note that
$|K^{i}_{\omega}-K^{j}_{\omega}+lb+mb^{\prime}|=|R^{T}(K^{i}_{\omega}-K^{j}_{\omega}+lb+mb^{\prime})|$ (132)
@@
$|K^{1}_{\omega}-K^{2}_{\omega}+lb+mb^{\prime}|\geq\frac{C_{0}}{|m|^{\tau}}$ (133)
@@
$|K^{1}_{\omega}-K^{2}_{\omega}+lb+mb^{\prime}|\geq\frac{2\pi}{\sqrt{3}}\left(x_{\omega}(\theta)+l_{1}+m\cdot f_{\omega}(\theta)\right)$ (134)
@@
Therefore, if $\theta\in\mathcal{D}(x_{\omega},f_{\omega})$, then (133) holds with $C_{0}=\frac{2\pi}{\sqrt{3}}C_{1}$. To show that this set has a large measure, we check the assumptions of Lemma B.1, as we did above. Again, we check the assumptions at $\alpha=0$, and argue by continuity.
We check (108). We choose $C_{3}$ in such a way that $R_{1}<\sqrt{2}$. In Figure 5 (left), we see that $\frac{|x_{\omega}^{\prime}(\theta)|}{|f_{\omega}^{\prime}(\theta)|}<1$ for all $\theta\in[0,2\pi)$, so we choose $C_{3}$ such that $C_{3}<1-\frac{|x^{\prime}_{\omega}(\theta)|}{|f^{\prime}_{\omega}(\theta)|}$ (recall that, since $\beta=\frac{\pi}{4}$, $\sqrt{2}\cos(\beta)=1$). Thus, to verify (108), it suffices to check the cases for which $|k|=1$, that is, $k=(\pm 1,0)$ and $k=(0,\pm 1)$. And, again choosing $C_{3}$ to be small enough, it suffices to check that $x^{\prime}_{\omega}(\theta)+k\cdot f^{\prime}_{\omega}(\theta)\neq 0$. Using a symbolic calculator, we find that this holds as long as
$\theta\not\in\mathcal{B}_{1}:=\left\{0,\frac{\pi}{3},\arctan(3\sqrt{3}),\pi-\arctan(3\sqrt{3}),\frac{2\pi}{3},\pi,\frac{4\pi}{3},\pi+\arctan(3\sqrt{3}),2\pi-\arctan(3\sqrt{3}),\frac{5\pi}{3}\right\}.$ (136)
(Figure 6 shows the plot of $x^{\prime}_{\omega}(\theta)+k\cdot f^{\prime}_{\omega}(\theta)$.)
Figure 5: $\frac{|x^{\prime}_{\omega}(\theta)|}{|f^{\prime}_{\omega}(\theta)|}$ and $|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}|/|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}|$ (blue: $\omega=+$, red: $\omega=$-, color online). The green lines represent the value 1. The plot on the left is used to verify (108): as we see, $\frac{|x_{\omega}^{\prime}(\theta)|}{|f_{\omega}^{\prime}(\theta)|}<1$, so, to verify (108), it suffices to check the cases for which $|k|=1$. The plot on the right is used to verify (110): as we see, $|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}|/|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}(\theta)}{|f^{\prime}_{\omega}(\theta)|}|<1$, so to verify (110), it suffices to check the cases for which $|k|=1$.
Figure 6: Left: $|x^{\prime}\pm f_{1}^{\prime}|$, right: $|x^{\prime}\pm f_{2}^{\prime}|$, in each graph, blue and green are $\omega=+$, red and yellow are $\omega=$-, blue and red are $\pm=+$, green and yellow are $\pm=$-. These plots are used to verify (108) for $|k|=1$. As we see, $x^{\prime}+k\partial f$ only reaches 0 at a finite number of points, enumerated in (136).
Finally, we check (110). We choose $\eta$ in such a way that $R_{2}<\sqrt{2}$. In Figure 5 (right), we see that $\left|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\right|<\left|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\right|$ for all $\theta\in[0,2\pi)$, so we choose $\eta$ such that $\eta<\sqrt{2}-\sqrt{2}\left|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\right|/\left|\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\right|$ (recall that, since $\beta=\frac{\pi}{4}$, $\cos(\frac{\pi}{2}-\beta)=\frac{1}{\sqrt{2}}$). Thus, to verify (110), it suffices to check the cases for which $|k|=1$, that is, $k=(\pm 1,0)$ and $k=(0,\pm 1)$. And, again choosing $\eta$ to be small enough, it suffices to check that $\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}+k\cdot\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\neq 0$. Using a symbolic calculator, we find that this holds as long as
[…]
(Figure 7 shows the plot of $\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}+k\cdot\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}$.)
Figure 7: Left: $|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\pm\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega,1}}{|f^{\prime}_{\omega}|}|$, right: $|\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}\pm\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega,2}}{|f^{\prime}_{\omega}|}|$, in each graph, blue and green are $\omega=+$, red and yellow are $\omega=$-, blue and red are $\pm=+$, green and yellow are $\pm=$-. These plots are used to verify (110) for $|k|=1$. As we see, $\frac{\partial}{\partial\theta}\frac{x^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}+k\cdot\frac{\partial}{\partial\theta}\frac{f^{\prime}_{\omega}}{|f^{\prime}_{\omega}|}$ only reaches 0 at a finite number of points, enumerated in (137).
@@
All in all, we have found that if we restrict the values of $\theta$ to a small interval $\theta_{0},\theta_{1}]$ that steers clear of any of the angles in $\mathcal{B}_{1}\cup\mathcal{B}_{2}$ (see (136) and (137)), and if we further restrict $\theta$ to an intersection of Diophantine sets:
$\theta\in[\theta_{0},\theta_{1}]\cap\bigcap_{\omega=\pm}\bigcap_{\sigma=\pm}\mathcal{D}(0,f_{\omega}(\sigma\theta))\cap\bigcap_{\omega=\pm}\bigcap_{\sigma=\pm}\mathcal{D}(x_{\omega}(\sigma\theta),f_{\omega}(\sigma\theta))$ (138)
@@
then (4) is satisfied for any value of $i,i^{\prime}$, $\omega,\omega^{\prime}$, and $y$ with a constant $C_{0}=O((\theta_{1}-\theta_{0})^{2})$. Because each set has an arbitrarily large measure (relative to $\theta_{0},\theta_{1}]$), their intersection also does.
@@ Appendix C Naive perturbation theory
The 2-point function function is computed using perturbation theory: formally,
@@ D.1 Symmetries of the system
$H_{1}$ and $H_{2}$ are invariant under (145) for the same reason as the conjugation and particle-hole symmetries. For the interlayer hopping, it is easiest to use the expression of $V$ in (16), which involves two integrals: one over $\widehat{\mathcal{L}}_{1}$ and one over $\widehat{\mathcal{L}}_{2}$. Let us discuss the invariance of the integral over $\widehat{\mathcal{L}}_{1}$, as the invariance of the other follows from a similar argument. We change variables in the integral: $k\mapsto-k+b_{1}+b_{2}$, as well as in the sum over $l$: $(l_{1},l_{2})\mapsto-(l_{1}+1,l_{2}+1)$, and in the sum over $\alpha$: $\alpha\mapsto\bar{\alpha}$. This changes $\tau_{l,\alpha}^{(1)}(k+lb)$ to $\tau_{-l-(1,1),\bar{\alpha}}^{(1)}(-k-lb)$. Now, by (17),